The current price of Estelle Corporation stock is $29.00. In each of the next two​ years, this stock price will either go up by 21% or go down by 21%. The stock pays no dividends. The​ one-year risk-free interest rate is 6.5% and will remain constant. Using the Binomial​ Model, calculate the price of a​ one-year put option on Estelle stock with a strike price of $29.00
Answer #1

Using the Binomial Model; Payoff table of Put option with strike price of $29:

Price after one year

Value of put option [Strike price ($29) – price]

Move up (u=35.09/29= 1.21)

$29 *1.21 = $35.09

$0 (price at the strike price so option will not exercise)

Current Price of stock($29)

Move down (d=22.91/29 =0.79)

$29 *(1-0.21) = $22.91

$6.09

Probability of moving up, P = e ^r*t – d / (u-d)

Where

Risk free rate, r = 6.5% per year or 0.065

Time period, t = 1 year

Factor of moving up, u = 1.21

Factor of moving down, d = 0.79

Therefore,

P = [e^ (0.065*1) – 0.79] / (1.21 -0.79)

P = 0.6599

And (1- P) = 1- 0.6599 = 0.3401 (probability of moving down)

Now the expected value of put option

= e^ - (0.065*1) * (1-P) * payoff of put option

= 0.9371 *0.3401 *$6.09

= $1.94

Therefore the value of the put option is $1.94